# Towards a Deeper Understanding of Reasoning Capabilities in Large Language Models

### v1

@@ Abstract
While large language models demonstrate impressive performance on static benchmarks, the true potential of large language models as self-learning and reasoning agents in dynamic environments remains unclear. This study systematically evaluates the efficacy of self-reflection, heuristic mutation, and planning as prompting techniques to test the adaptive capabilities of agents. We conduct experiments with various open-source language models in dynamic environments and find that larger models generally outperform smaller ones, but that strategic prompting can close this performance gap. Second, a too-long prompt can negatively impact smaller models on basic reactive tasks, while larger models show more robust behaviour. Third, advanced prompting techniques primarily benefit smaller models on complex games, but offer less improvement for already high-performing large language models. Yet, we find that advanced reasoning methods yield highly variable outcomes: while capable of significantly improving performance when reasoning and decision-making align, they also introduce instability and can lead to big performance drops. Compared to human performance, our findings reveal little evidence of true emergent reasoning. Instead, large language model performance exhibits persistent limitations in crucial areas such as planning, reasoning, and spatial coordination, suggesting that current-generation large language models still suffer fundamental shortcomings that may not be fully overcome through self-reflective prompting alone. Reasoning is a multi-faceted task, and while reasoning methods like Chain of thought improves multi-step reasoning on math word problems, our findings using dynamic benchmarks highlight important shortcomings in general reasoning capabilities, indicating a need to move beyond static benchmarks to capture the complexity of reasoning.
[…]
123
1 Introduction
@@
A key objective of artificial intelligence has been the development of intelligent agents that can perceive the environment and make autonomous decisions [31]. The emergence of large language models (LLMs) has significantly advanced this objective, showing strong capabilities in the solution of various natural language processing tasks, such as solving mathematical problems, coding, reading comprehension, translation, summarizing and answering questions [1, 4, 20]. These findings show a promising direction towards autonomous agents, and recent studies have begun to explore methods for enabling agents to learn dynamically. For instance, investigations into self-reflection mechanisms allow agents to evaluate their past actions and refine future strategies [22], while work on iterative prompting and environmental feedback aims to facilitate continuous learning from experience [35]. LLMs have also been investigated to serve as a controller to solve environmental tasks through reasoning and planning [33, 36, 19, 13]. The ability of LLM agents to learn and adapt in dynamic environments has yet to be definitively proven. While these models excel at in-context learning—generalizing from minimal examples—their reliance on statistical prediction and lack of long-term memory often limit their effectiveness in dynamic settings [25]. Furthermore, achieving optimal performance on specialised tasks typically requires either fine-tuning with human-annotated data [16] or reliance on careful prompt engineering [21]. These processes are resource-intensive, reducing the flexibility of deploying LLMs in real-world, constantly evolving applications. This study aims to understand the following question: "To what extent can LLM agents autonomously learn and adapt to novels tasks in dynamic environments?" Specifically, we investigate whether in-context mechanisms can improve continuous learning and multi-step reasoning across various challenge levels. Our contributions are as follows  [^1] Code available at https://github.com/ann-w/towards-a-deeper-understanding-of-reasoning-capabilities-in-large-language-models :
@@
  • We present a systematic comparison of open-souce LLMs on dynamic decision-making tasks. We quantify three prompting strategies—reflection, heuristic mutation, and planning—across SmartPlay [32]: a benchmark to evaluate capabilities of intelligent agents. For simple reactive task, we find that excessive reasoning harm performance for smaller models, while larger models are more robust. Larger models generally achieve higher scores in line with scaling laws [7], but carefully designed prompts can let small models match or surpass the baseline performance of larger models. Advanced prompting techniques primarily benefit smaller models on complex planning and reasoning tasks, and offer less value to already high-performing large language models.
@@ 2 Background
Recent studies have increasingly focused on LLM capabilities as agents in complex, interactive text-based games. These environments present unique challenges requiring puzzle-solving and planning that stretch beyond the inherent statistical prediction capabilities of LLMs. In response, a spectrum of prompt-level techniques has emerged to improve LLM reasoning capabilities without the need for fine-tuning of the parameters. Chain-of-Thought (CoT) encourages the decomposition of complex tasks into sequential steps [30], while Self-Refine leverages iterative feedback to improve output quality [14]. ReAct [35] and Reflexion [36] integrate reasoning with action and self-reflection, enabling agents to continuously assess and adjust their strategies as they interact with the environment.
Others emphasise iterative plan refinement: AutoPlan [17] revises a natural-language plan while solving household tasks and multi-hop questions; DEPS [29] iteratively refines plans by summarizing situations, diagnosing errors, and selecting sub-goals in open worlds; RCI [8] refines plans on computer tasks. Others predict future states and rewards from the interaction history to steer their plan in short-horizon puzzles [12].
Finally, inspired by the successes of evolutionary approaches in other domains, researchers have begun to apply evolutionary strategies to self-optimise LLM performance. EvoPrompt [6] uses genetic algorithms and differential evolution to iteratively refine prompts, while metaheuristic methods in frameworks like PLUM [18] and LLaMEA [27] generate, mutate, and select effective algorithms based on performance feedback. This line of research underscores the potential for evolutionary algorithms to minimize manual tuning.
@@
Despite the proposal of diverse prompt techniques, methods have usually been studied in isolation. This study offers a more comprehensive evaluation by systematically testing different prompting strategies on the SmartPlay suite to test agent capabilities on nine different challenges, making it a more comprehensive benchmark than most studies that focus on a single domain, such as text-based household tasks [23] or grid-world instructions [3].
Figure 1: High-level representation of the agent framework. At each timestep, the agent takes an action, receives a reward and moves to the next state. The agent can receive additional guidance from three modules. Reflection performs retrospective analysis based on the past trajectory per timestep. The Oracle evolves heuristics via mutation to uncover environment dynamics based on past reflections and trajectory after each episode. The Planner simulates future states and rewards to recommend the best action based on the past trajectory and current reflection per timestep.
3 Methodology
Our framework is presented in Figure 1. At each timestep, the agent interacts with the environment by taking an action, receives a reward and moves to the next state. Reflection performs a retrospective analysis of the agent’s trajectory at each timestep. Oracle evolves heuristics after each episode by mutating them based on past reflections and trajectory to capture the generic episode dynamics. The Planner simulates future states and the cumulative reward based on the trajectory and reflection. The prompts for all modules described below can be found in the Supplementary Material.
@@ 3 Methodology
For our experiments, we use open source LLMs, including Llama3-8B, Mistral-Nemo-12b, DeepSeek-R1-14b and Llama3.3-70B to investigate how size and architecture influence performance. We ran experiments for the largest model, Llama-3.3-70B, on two A100 machines. The smaller models were run on NVIDIA RTX 4090, RTX 3060, or A40 GPUs. We cover each module in more detail below.
3.1 Agent
At each time step $t$, the agent selects an action $a_{t}$ to maximize the sum of rewards across episodes: $\max_{\pi}\sum_{k=1}^{T}r_{k}$. Using only the information within the prompt (game manual, task, history, current state, possible actions), the agent adjusts its policy $\pi$.
We evaluate the agent where the following module combinations are enabled: In the first strategy, only Reflection is enabled, where at each timestep, the agent receives a reflection on its past actions in the current episode, and the agent is prompted to take the next best action. In the second strategy, Reflection and Oracle are enabled, where the agent receives both heuristics (generated by the Oracle and kept consistent within an episode but mutated between episodes) and the aforementioned reflection over past episode trajectory per timestep, and the agent is prompted to take the next best action. In the third strategy, Reflection and Planner are enabled, where at each timestep the agent receives a reflection, and the Planner suggests a subsequent action, which the agent executes.
3.2 Reflection
Inspired by Reflexion [36], we add a reflection module so that the agent can reflect on the past trajectory within each episode. The reflection module reviews the agent’s past sequence of state, action, reward, and next state. This process involves comparing the sequence of actions taken against the game’s objectives and identifying adjustments that could lead to higher cumulative rewards. Reflections are generated at each time step within an episode, they are reset between episodes because the game’s dynamics may change.
3.3 Oracle
The Oracle generate heuristics from past reflections to optimize the agent’s policy, continuously refining these rules through a (1+1) evolutionary strategy over successive episodes. This process minimizes the need for manual prompt engineering while enabling the agent to adapt to a dynamic environment. The Oracle’s task is to identify general patterns across episodes; consequently, the generated heuristics remain consistent within a single episode. Specifically, in the first episode, the agent explores the environment, and the Oracle uses the resulting trajectory and reflections to create initial heuristics (the parent). Following each subsequent episode, the Oracle generates and evaluates a single offspring by mutating the parent heuristics. If the offspring’s performance is better than the parent’s, the offspring replaces the parent; otherwise, the parent heuristics are retained. This "mutation" process enables the LLM to iteratively refine the heuristics by adding, modifying, or removing rules to better align with observed data.
3.4 Planner
While the Reflection and Oracle modules analyze past trajectories and feedback to refine the agent’s strategy, the Planner is forward-looking. The planner’s objective is to simulate possible action sequences up to three steps ahead and select the action with the highest expected cumulative reward, based on the game manual, objectives, trajectory, reflection and current observation. We explored but did not proceed with adding heuristics as input to the planner because it led to a deviation from the response format, possibly because the instructions became obscured within the lengthy context.
3.5 Memory
The Memory maintains both episodic and persistent data. The episodic memory tracks the current episode’s trajectory. The persistent memory stores the best heuristic to be mutated across episodes. The episodic memory is limited to the trajectory of the current episode due to the varying dynamics that can occur within each episode.
4 Results
Figure 2: Heatmap illustrating the relative importance of skills for success in each environment (numbers from SmartPlay [32] and normalized). For Bandit, understanding odds and learning from past data is important. Rock, Paper, Scissors requires instruction following, opponent-based learning, and long-text comprehension. Hanoi emphasizes planning, reasoning, and careful instruction following. Messenger requires long-text interpretation, spatial reasoning, probabilistic thinking, and generalization.
4.1 Test Environments
@@
We evaluate our framework on four Smartplay environments [^2] https://github.com/microsoft/SmartPlay. The challenges are shown in Figure 2. In Bandit, the agent must balance exploration and exploitation to discover which bandit yields higher payoff. Rock Paper Scissors (RPS) tests probabilistic reasoning against an opponent whose move distribution is biased and randomly shuffled. In Tower of Hanoi, the agent must plan and reason spatially to move three disks across rods, with the rule that a larger disk can never rest on a smaller one. Messenger challenges the agent’s ability to understand text with synonyms and to use that understanding to move around, avoid an enemy, and deliver a message to the goal. Task horizons and episode counts follow Wu et al. [32] (See also Supplementary Material), except for Messenger, where the rollout is extended from 4 to 10 steps, because the initial placement of the enemy and goal made it often impossible to complete the task within the original four steps. We also reduced training to 20 episodes after observing no further performance improvements beyond that point.
@@
Table 1: The table shows the minimum, median and maximum average scores over three runs. The top row shows the human baseline for each game, taken from [32]. Darker colours indicate closer performance to human baseline. For Bandit, the score improves almost linearly with model size. For simple reactive tasks, excessive reasoning harms small models’ performance. Rock Paper Scissors demands adaptation to opponent behaviour in which model size is the main driver of better performance. Reflection + Oracle and Reflection + Planner benefits mostly smaller models in the more complicated Hanoi and Messenger tasks. Yet we observe large variability in runs. Note that the human baseline is not directly comparable with our results, because it was obtained under a 4-step, 100-episode setting, whereas our evaluation uses a 10-step horizon and 20 episodes.
@@ 4.2 Results and Analysis
Table 1 shows the minimum, median and maximum average scores over 3 runs for each strategy. [^3] All results are averaged over three independent runs; the budget is capped at three because of the computational cost imposed by LLaMA3.3-70B. Scoring varies by environment: for Bandit and RPS, it’s the number of optimal actions; for Hanoi, it’s the count of successfully moved discs; and for Messenger, score equals reward (with -1 for errors and +1 for successful message actions). Human baselines are taken from [32]. A ranking with statistical tests of all model-strategy pairs can be found in the Supplementary Material. In the next sections, we report results as median score [min, max].
4.2.1 Model Size and Performance
Larger parameter counts generally achieve higher scores (see Table 1). For instance, on Bandit, Llama3.3-70b $41.70\,[40.90-41.90]$ outperforms deepseek-r1-14b $41.00\,[40.55-41.40]$, Mistral-Nemo-12b $34.20\,[30.95-36.30]$ and Llama3-8b $40.35\,[37.45-41.65]$. Model performance increases with size across all game baselines. This performance gap widens as the games become more challenging. For example, Llama3.3-70b achieves a score of 2.00[2.002.00]2.00delimited-[]2.002.002.00\,[2.00-2.00]2.00 [ 2.00 - 2.00 ] in Hanoi, significantly outperforming Llama3-8b’s median score of $0.20\,[0.20-0.30]$. Overall, Llama3.3-70b leads, closely followed by deepseek-r1-14b. Only Llama3-8b attains higher baseline median scores than the larger Mistral-Nemo-12b.
Prompting techniques can close the gap between smaller and larger models on specific tasks. For example, Llama3-8b with Reflection + Oracle $26.00\,[12.00-36.00]$ outperforms Llama3.3-70b’s baseline of $22.20\,[21.55-25.05]$ on RPS. In Messenger, Mistral-Nemo-12b with Reflection + Planner 1.00[1.001.00]1.00delimited-[]1.001.001.00\,[-1.00-1.00]1.00 [ - 1.00 - 1.00 ] exceeds Llama3.3-70b’s baseline 0.10[0.100.10]0.10delimited-[]0.100.100.10\,[-0.10-0.10]0.10 [ - 0.10 - 0.10 ], as does DeepSeek-R1-14B with Reflection + Oracle $1.00\,[0.00,1.00]$ and Reflection + Planner $1.00\,[0.40-1.00]$, matching Llama3.3-70b’s maximum score. However, these improvements are inconsistent. For instance, Mistral-Nemo-12b’s Reflection + Planner score in RPS ranges from 10.00 to 33.00, and DeepSeek-R1-14B’s Reflection + Oracle ranges from 0.00 to 1.00, indicating unreliable success.
In conclusion, model size has the highest impact on overall performance. While prompting strategies can enable mid-sized models to match or close the gap with large models, their effectiveness is inconsistent, and the performance gain by increased model parameters is difficult to overcome by in-context learning alone.
Figure 3: Heat-map of Pearson correlation coefficients between each strategy’s performance gain over the Base prompt and challenge dimensions over three evaluation runs per environment. DeepSeek-R1-14B benefits most on Instruction-Following ($r=0.87\!-\!0.94$) and Long-Text Understanding ($r=0.69\!-\!0.79$), while slightly harming Generalisation and Understanding the Odds. Llama3-8B shows its largest lift on Instruction Following ($r=0.93\!-\!0.94$) and on Learning from Interactions when an Oracle is present ($r=0.73$). For Llama3.3-70B, all strategies yield near-perfect alignment with Learning from Interactions and gains for Understanding the Odds, Generalization and Instruction Following. Mistral-Nemo-12B consistently improves Instruction-Following ($r=0.82\!-\!0.98$) and moderately boost Long-Text Understanding ($r=0.55\!-\!0.81$), with little impact on other challenges. Overall, strategies improve Instruction Following across all models. For smaller and mid-sized models, these strategies also contribute to better long-text understanding, a benefit less pronounced in the already capable Llama3.3-70B. Conversely, for the larger Llama3.3-70B, the strategies improve generalization, an area where smaller models still face challenges. While no specific strategy type strongly correlates with a particular challenge, the application of strategies is associated with performance gains in specific challenge types.
4.2.2 Strategies versus Challenge Type
[…]
Overall, strategies improve Instruction Following across all models. For smaller and mid-sized models, these strategies also contribute to better long-text understanding, a benefit less pronounced in the already capable Llama3.3-70B. Conversely, for the larger Llama3.3-70B, the strategies improve generalization, an area where smaller models still face challenges.
Next, we provide an analysis of the results per environment.
4.2.3 TwoArmedBandit

### v2

# Reasoning Capabilities of Large Language Models on Dynamic Tasks

@@ Abstract
Large language models excel on static benchmarks, but their ability as self-learning agents in dynamic environments remains unclear. We evaluate three prompting strategies: self-reflection, heuristic mutation, and planning across dynamic tasks with open-source models. We find that larger models generally outperform smaller ones, but that strategic prompting can close this performance gap. Second, an overly long prompt can negatively impact smaller models on basic reactive tasks, while larger models show more robust behaviour. Third, advanced prompting techniques primarily benefit smaller models on complex games, but offer less improvement for already high-performing large language models. Yet, we find that advanced reasoning methods yield highly variable outcomes: while capable of significantly improving performance when reasoning and decision-making align, they also introduce instability and can lead to big performance drops. Compared to human performance, our findings reveal little evidence of true emergent reasoning. Instead, large language model performance exhibits persistent limitations in areas like planning and spatial coordination, suggesting that large language models still suffer fundamental shortcomings that may not be fully overcome through self-reflective prompting alone. Reasoning is a multi-faceted task, and while methods like Chain-of-thought improves multi-step reasoning on math word problems, our findings using dynamic benchmarks highlight important shortcomings in general reasoning capabilities, indicating a need to move beyond static benchmarks to capture the complexity of reasoning.
Index Terms:
Large Language Models, In-context learning, Reasoning
Figure 1: Top: Overview of the four prompting strategies, with each strategy’s added text shown in bold within the prompt. At every timestep in the environment, the agent chooses its next move solely from the text in its prompt: game manual, history as (state, action, reward, next state) tuples from episode start to the current step, current observation, and the list of legal actions. The example prompts in the figure omit the manual and the action list for brevity, though it is always present at runtime. Bottom: The agent can receive additional guidance from three modules. Reflection performs retrospective analysis based on the past trajectory per timestep. The Oracle evolves heuristics via mutation to uncover environment dynamics based on past reflections and trajectory after each episode. The Planner simulates future states and rewards to recommend the best action based on the past trajectory and current reflection per timestep.
I Introduction
@@
A key objective of artificial intelligence has been the development of intelligent agents that can perceive the environment and make autonomous decisions. The emergence of large language models (LLMs) has significantly advanced this objective, showing strong capabilities in the solution of various natural language processing tasks, such as translation, reading comprehension, and answering questions [1]. These findings show a promising direction towards autonomous agents, and recent studies have begun to explore methods for enabling agents to learn dynamically. For instance, investigations into self-reflection mechanisms allow agents to evaluate their past actions and refine future strategies [2], while work on iterative prompting and environmental feedback aims to facilitate continuous learning [3]. LLMs have also been investigated to serve as a controller to solve environmental tasks through reasoning and planning [4]. The ability of LLM agents to learn in dynamic environments has yet to be definitively proven. While these models excel at in-context learning—generalizing from minimal examples—their reliance on statistical prediction and lack of long-term memory often limit their effectiveness in dynamic settings [5]. Achieving optimal performance on specialised tasks typically requires either fine-tuning with human-annotated data [6] or reliance on careful prompt engineering [7]. These processes are resource-intensive, reducing the flexibility of deploying LLMs in real-world, constantly evolving applications. Specifically, we investigate whether in-context mechanisms can improve continuous learning and multi-step reasoning across various challenges and aim to answer: To what extent can LLM agents autonomously learn and adapt to tasks in dynamic environments?”. Our contributions are as follows  [^1] Code available at https://github.com/ann-w/towards-a-deeper-understanding-of-reasoning-capabilities-in-large-language-models :
@@
  • We present a comparison of open-source LLMs (Llama3-8B, Mistral-Nemo-12b, DeepSeek-R1-14b and Llama3.3-70B) using three prompting strategies: reflection, heuristic mutation, and planning across SmartPlay [8]: a benchmark to evaluate agent capabilities in dynamic tasks. For simple tasks, excessive reasoning harms performance for smaller models, while large models are more robust. Although larger models typically lead in performance, carefully crafted prompts can push smaller models past large-model baselines. These advanced prompting techniques yield the greatest gains for smaller models on complex planning and reasoning tasks, while offering only modest improvements to already high-performing models.
@@ II Background
Recent studies have increasingly focused on LLM capabilities as agents in complex, interactive text-based games. These environments present unique challenges requiring puzzle-solving and planning that stretch beyond the inherent statistical prediction capabilities of LLMs. In response, a spectrum of prompt-level techniques has emerged to improve LLM reasoning capabilities without the need for fine-tuning of the parameters. Chain-of-Thought (CoT) encourages the decomposition of complex tasks into sequential steps [9], while Self-Refine leverages iterative feedback to improve output quality [10]. ReAct [3] and Reflexion [11] integrate reasoning with action and self-reflection, enabling agents to continuously assess and adjust their strategies as they interact with the environment. Others emphasise iterative plan refinement: AutoPlan [12] revises a natural‑language plan while solving household tasks and multi‑hop questions; DEPS [13] iteratively refines plans by summarizing situations, diagnosing errors, and selecting sub-goals in open worlds. Others predict future states and rewards from the interaction history to steer their plan in short‑horizon puzzles [14]. Researchers have also begun to apply evolutionary strategies to self-optimise LLM performance. EvoPrompt [15] uses genetic algorithms and differential evolution to iteratively refine prompts, while metaheuristic methods in frameworks like LLaMEA [16] generate, mutate, and select effective algorithms based on performance feedback. This line of research underscores the potential for evolutionary algorithms to minimize manual tuning.
@@
Despite the proposal of diverse prompt techniques, methods have usually been studied in isolation. This study offers a comprehensive evaluation by systematically testing different prompting strategies on the SmartPlay suite to test agent capabilities on nine different challenges, making it a more comprehensive benchmark than most studies focusing on a single domain.
III Methodology
III-A Agent Framework
Our agent operates in discrete episodes, each consisting of a sequence of timesteps from a start state to a terminal state. At each timestep $t$, the agent observes the current state and selects an action $a_{t}$, then receives a reward $r_{t}$ and transitions to the next state. The objective within a single episode (of horizon $T$ steps) is to maximize the cumulative reward $R=\sum_{k=1}^{T}r_{k}$, where $T$ is the episode length and $r_{k}$ is the reward at step $k$. We implement the agent’s policy $\pi$ using a LLM that decides the next action based solely on the information in its prompt at each step. This prompt includes the game’s manual, the task objective, the episode history recorded as a sequence of ⟨state, action, reward, next-state⟩ tuples, the current observation, and the list of possible actions. No parameter fine-tuning is performed; the agent relies on in-context reasoning to choose actions that yield high rewards.
The base agent can be augmented with three modules: Reflection, Oracle, and Planner (Figure 1). Reflection performs a retrospective analysis of the agent’s trajectory at each timestep, providing feedback on recent actions. Oracle operates at the episode level, evolving heuristic rules between episodes to help the agent adapt to changing environment dynamics. Planner is a forward-looking component that simulates future action sequences to recommend the best next action at each timestep. These modules interact with the agent as follows: after the agent takes an action and observes the outcome, the Reflection module (if enabled) generates feedback which is added to the agent’s context; when the Planner is enabled, the agent uses it before deciding an action to explore hypothetical scenarios; and at the end of each episode, if the Oracle is enabled, it updates or mutates a set of heuristics that will be provided to the agent in the next episode.
III-A1 Reflection
Adapted from Reflexion [11], this module enables the agent to analyze its own behavior within the current episode and adjust accordingly. After each action–state transition, the agent inspects the trajectory (stored in memory), compares it to the task objective, and produces a brief suggestion. This reflection can be seen as a meta-level feedback: it compares the agent’s actions against the game’s goals and highlights potential mistakes or better strategies that could lead to a higher reward. We reset the reflection at the start of each new episode, since feedback from a past episode may not be directly applicable when the environment’s conditions change in the next episode. For instance, the profitable bandit might be Bandit 1 in the first episode, but switch to Bandit 2 in the second episode.
III-A2 Oracle
The Oracle module aims to capture cross-episode learning by evolving heuristic rules that guide the agent in future episodes using a simple (1+1) evolutionary strategy applied to textual heuristics: each episode generates one mutated offspring heuristic that replaces the parent only if it achieves a higher reward. After the first episode, the Oracle analyzes the trajectory and reflections from that episode to formulate an initial set of heuristic guidelines. These heuristics are intended to generalize beyond a single episode. For example, a rule might be “Always try the lever that gave a high reward previously” in a multi-armed bandit game. In subsequent episodes, the Oracle continuously refines these rules: at the end of each episode, it generates a mutated version of the current heuristic set (an offspring) by adding, removing, or modifying one of the rules. We then test the offspring in the next episode. If the agent achieves a higher cumulative reward with the new heuristics, the mutated heuristics are kept as the new “parent” set and stored in memory; if not, the Oracle reverts to the previous best heuristics. This evolutionary process allows the agent to adapt to dynamic environments without manual prompt tuning. The Oracle’s heuristics remain fixed during each episode (providing stable guidance within that episode), and only change between episodes as the agent learns from new experience.
III-A3 Planner
While Reflection and Oracle focus on the past, the Planner looks ahead to improve decision-making. The Planner simulates possible future trajectories from the current state at each timestep, effectively doing a short-horizon lookahead. Using the game manual, current state, and latest reflection, the Planner simulates rollouts up to three steps, estimates their cumulative rewards, and returns the immediate action with the highest expected return. The Planner’s recommendation for the next action is followed by the agent when the Planner strategy is enabled. While we also experimented with providing Oracle’s heuristics to the Planner, we found that the long context caused the LLM to deviate from the prescribed answer format.
III-B Models and Hardware Configuration
@@
We use open source LLMs, including Llama3-8B, Mistral-Nemo-12b, DeepSeek-R1-14b and Llama3.3-70B to investigate how size and architecture affect performance. We ran experiments for the largest model, Llama-3.3-70B, on two A100 machines. Smaller models were run on NVIDIA RTX 4090, RTX 3060, or A40 GPUs.
IV Results
IV-A Test Environments
@@
We evaluate our framework on four Smartplay environments [^2] https://github.com/microsoft/SmartPlay. In Bandit, the agent must balance exploration and exploitation to discover which bandit yields higher payoff. Rock Paper Scissors (RPS) tests probabilistic reasoning against an opponent whose move distribution is biased and randomly shuffled. In Tower of Hanoi, the agent must plan and reason spatially to move three disks across rods, with the rule that a larger disk can never rest on a smaller one. Messenger challenges the agent’s ability to understand synonyms and to use that understanding to move around, avoid an enemy, and deliver a message to the goal. We adopt episode counts of [8], except that Messenger rollouts are lengthened from 4 to 10 steps because the original horizon often made the goal unreachable, and training is capped at 20 episodes after observing no further performance improvements beyond that point.
@@
TABLE I: The table shows the minimum, median and maximum average scores over three runs. The top row shows the human baseline for each game, taken from [8]. Darker colours indicate closer performance to human baseline. For Bandit, the score improves almost linearly with model size. For simple reactive tasks, excessive reasoning harms small models’ performance. Rock Paper Scissors demands adaptation to opponent behaviour in which model size is the main driver of better performance. Reflection + Oracle and Reflection + Planner benefits mostly smaller models in the more complicated Hanoi and Messenger tasks. Yet we observe large variability in runs. Note that the human baseline is not directly comparable with our results, because it was obtained under a 4-step, 100-episode setting, whereas our evaluation uses a 10-step horizon and 20 episodes.
@@ IV-B Results and Analysis
Table I shows the scores for each strategy. For each combination of strategy and game environment, we ran three independent runs [^3] All results are averaged over three independent runs; the budget is capped at three because of the computational cost imposed by LLaMA3.3-70B.. In each run, the agent with a given strategy played through the fixed number of episodes, and we computed the average score across those episodes for that run. Then, we rank these three average scores. The highest score is reported as the maximum, the lowest score as the minimum, and the middle score as the median. Each environment defines success differently: in Bandit and Rock–Paper–Scissors, the score represents the frequency of optimal actions; in Hanoi, it is the count of moved disks to the goal and in Messenger, it is the accumulated reward, where correct message actions yield +1 and invalid moves incur -1. Human baselines are taken from [8]. In the following sections, we report results as median score [min, max].
IV-B1 Model Size and Performance
Larger parameter counts generally achieve higher scores (see Table I). For instance, on Bandit, Llama3.3-70b 41.70[40.9041.90]41.70\,[40.90-41.90] outperforms deepseek-r1-14b 41.00[40.5541.40]41.00\,[40.55-41.40], Mistral-Nemo-12b 34.20[30.9536.30]34.20\,[30.95-36.30] and Llama3-8b 40.35[37.4541.65]40.35\,[37.45-41.65]. Model performance increases with size across all game baselines. This performance gap widens as the games become more challenging. Prompting techniques can close this gap on specific tasks. For example, Llama3-8b with Reflection + Oracle 26.00[12.0036.00]26.00\,[12.00-36.00] outperforms Llama3.3-70b’s baseline of 22.20[21.5525.05]22.20\,[21.55-25.05] on RPS. In Messenger, Mistral-Nemo-12b with Reflection + Planner 1.00[1.001.00]1.00\,[-1.00-1.00] exceeds Llama3.3-70b’s baseline 0.10[0.100.10]0.10\,[-0.10-0.10]. However, these improvements also increase score variance across all models. For instance, Mistral-Nemo-12b’s Reflection + Planner score in RPS ranges from 10.00 to 33.00, and DeepSeek-R1-14B’s Reflection + Planner ranges from 0.00 to 2.00 in Hanoi. This indicates that advanced prompting strategies can yield significant improvements but are also unstable and can lead to big performance drops. Specifically, adding the Planner module on top of Reflection usually widens the variance more than the Reflection + Oracle strategy. While prompting strategies can offset smaller parameter counts, larger models still offer the most consistent advantage. Achieving performance improvements solely through in-context learning is challenging and often comes with increased variability for all model sizes.
Figure 2: Normalized performance difference compared to baseline for various models and prompting strategies across nine challenges. Llama3-8B and Mistral-Nemo-12B generally benefit, especially from Reflection + Planner, with consistent improvements. In contrast, DeepSeek-R1-14B shows minimal gains, and Reflection can even degrade its performance. While some tasks improve slightly for Llama3.3-70B, many tasks show regressions (e.g., Planning score reduction with Reflection + Planner). All models consistently improve on Instruction Following. However, the overall impact of advanced prompting varies with model capacity and task, requiring tailored strategies for each model and challenge type.
IV-B2 Strategies versus Challenge Type
Figure 2 shows the difference each prompting strategy achieves relative to the baseline on the nine SmartPlay skills. For every game $g$ we min–max scale the raw mean score $r_{m,s,g}$ to $x_{m,s,g}\in[0,1]$ ($m=$ model, $s=$ strategy). The score difference is calculated as $\delta_{m,s,g}=x_{m,s,g}-x_{m,\text{Base},g}$. SmartPlay assigns weights $w_{g,d}\in\{0.33,0.67,1\}$ that indicate how strongly game $g$ tests challenge dimension $d$. Weighted aggregation then gives
$\Delta_{m,s,d}=\frac{\sum_{g}w_{g,d}\,\delta_{m,s,g}}{\sum_{g}w_{g,d}},$ (1)
where $d$ indexes the challenge dimensions; a positive (negative) $\Delta_{m,s,d}$ means strategy $s$ increases (decreases) model $m$’s performance on dimension $d$ relative to the baseline.
Llama3-8B benefits most from the Reflection + Planner strategy, achieving large gains (+0.19) on Planning and Reasoning challenges. The other two strategies also result in score increases compared to the baseline. Similarly, Mistral-Nemo-12B shows improvements across most challenges, with the addition of an Oracle or Planner yielding greater gains than reflection alone. DeepSeek-R1-14B shows minimal improvements with prompting strategies: Long Text Understanding shows a modest improvement of 0.05, and Reflection alone can even drag its performance down (for instance, –0.08 on Error Handling). For the largest model, Llama3.3-70B, the structured prompts produce mixed results: they offer small gains in some challenges, such as improving Learning from Interactions by +0.05 and Understanding the Odds by +0.04, but lead to adverse effects in others. For instance, the Reflection + Planner prompt reduces the Planning score by approximately 0.15 relative to the baseline. All models consistently improve on Instruction Following. However, the impact of prompting strategies varies significantly across different models. While smaller models show improvements across all nine challenges, larger models show a regression on several challenges. This suggests that advanced prompting strategies can interfere with a high-capacity model’s inherent abilities, highlighting the importance of tailoring the prompting strategy to the specific model and challenge. Next, we provide an analysis of the results per environment.
IV-B3 TwoArmedBandit